OpenStudy (anonymous):

The sum of three numbers is -48. The first number is twelve more than the second number, and the third number is 20 less than twice the second number. Find the three numbers.

OpenStudy (ranga):

Let second number be x. First number = x + 12 Third number = 2x - 20 Sum of all three numbers = -48 Add up the numbers: x + (x + 12) + (2x - 20) = -48 Solve for x.

OpenStudy (anonymous):

the x = 1?

OpenStudy (ranga):

No x is not 1. Show me your work and I can see where the mistake is.

OpenStudy (anonymous):

it says solve for x.. so i do x + (x + 12) + (2x - 20) = -48?

OpenStudy (ranga):

Yes. Simplify the left side.

OpenStudy (anonymous):

multiply??

OpenStudy (ranga):

No add. x + x + 12 + 2x - 20

OpenStudy (anonymous):

oh its adding.

OpenStudy (ranga):

Yes, what do you get?

OpenStudy (anonymous):

i was trying to multiply ok now i not lost thanks

OpenStudy (anonymous):

i got -16 ?

OpenStudy (anonymous):

i simplify -20 + - 12

OpenStudy (anonymous):

then i 2x = -32

OpenStudy (anonymous):

is this correct???

OpenStudy (ranga):

No. But if you do it step by step here I can spot any mistakes. Start with x + (x + 12) + (2x - 20) = -48 Simplify the left side first and show me just that step.

OpenStudy (anonymous):

OK X = 14

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

i add up the x = 4

OpenStudy (anonymous):

x=4

OpenStudy (ranga):

Sorry no. x + (x + 12) + (2x - 20) = -48 x + x + 2x + 12 - 20 = -48 4x - 8 = -48 add 8 to both sides 4x = -40 divide both sides by 4 x = -10 The three numbers are: x+12, x and 2x - 20. Put x = -10 (-10+12), -10, 2(-10) - 20 2, -10, -40 are the three numbers.

OpenStudy (anonymous):

OH I forgot about 4x thanks! i understand now

OpenStudy (anonymous):

i keep adding without the 4x

OpenStudy (ranga):

ok, cool.